OpenStudy (anonymous):

what is the intergration of 1 divide by x

Parth (parthkohli):

\[\int \dfrac{1}{x}dx = \ln(x)\]

Parth (parthkohli):

\[\ln(x) + C\]

OpenStudy (anonymous):

can you explain.

Parth (parthkohli):

\[\dfrac{d}{dx}\ln(x) = \dfrac{1}{x} \implies \int \dfrac{1}{x}dx = \ln(x) + C\]

OpenStudy (anonymous):

thnaks

OpenStudy (anonymous):

what is the intergration of In x divide x?

OpenStudy (anonymous):

\[(\ln x)^{2}/2 +c \] Substitute ln x = t and then integrate.